40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4694     Entry time: Wed May 11 22:52:55 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: status update and plan 

Rana forced me to write this entry for summary because he didn't come to the 40m meeting today.

Status update :

    Interferometer Input Beam alignment with the PZTs.

     60 % done. The rest of the 40 % is to make the procedures automated.

     The beam spots on ITMY and ETMY are centered within ~1 mm accuracy.

     PRM, BS, ITMX, & ITMY actuator calibrations and PRC/MICH error calibrations

    Ongoing: First we will do it by hand, then some scripts will be made for the calibraion and resultant noise budget.

     F2A Suspension filter calculations.

      ETMY and ITMY are done. Need volunteer for ETMX, ITMX, BS, PRM, & SRM !!

      Bounce-Roll notch filters

      Leo is working on it. 25% complete...

     DC signal from RFPDs

      The RFPDs have a local SMA DC output as well as a DC output from their PD Interface cards in the LSC rack. We have hooked up some of the PD Interface DC outputs to the LSC ADCs but not tested??

Next steps:

  Installation of a temporary (Thorlabs) DCPD on either POY to see the intra-cavity power in the PRC. It would be ridiculous to put detectors on POX or POP since they're still clipped.

  D-phase and amplitude imbalance adjustment of the demod baords.

          make a script which uses pynds.

  Alignment of the full interferomter, starting from the X arm

  Loss measurements for the arms

  Schnupp asymmetry measurement

  

ELOG V3.1.3-